You are helping a manufacturing firm decide whether is should invest in a new plant. The initial investment is expected to be $50 million, and the plant is expected to generate after-tax cashflows of $5 million a year for the next 20 years. There will be an additional investment of $20 million needed to upgrade the plant in 10 years. If the discount rate is 10%. 

 

a. Estimate the Net present Value of the project.

b. Prepare a Net Present Value Profile for this project.

c. Estimate the Internal Rate of Return for this project. Is there any aspect of the cashflows that may prove to be a problem for calculating IRR?



1
Expert's answer
2021-02-24T06:32:25-0500

1Initial investment=50000000

2Cash Inflow per year for 20 years=5000000

3Investment after 10 years=20000000

4Discount Rate r=10%

5Net present value of the project= (- Initial Investments+ Net Present Value of all cash inflow- present value of investment after 10 years)

6Initial Investments-5,00,00,000.00

7Net Present Value of all cash inflow = PV(0.1,20,5000000)4,25,67,818.60

8Present value of investment after 10 years= -20000000/(1.1^10)-7710865.789

9Net present value of the project=-1,51,43,047.19


ii) The project APPARENTLY may not be accepted due to its negative  Net present value.

But considering positive IRR as achieved below, it may be considered to be acceptable.


ii) IRR can be calculated by drawing cash flow chart for the respective active project year.

The same can be depicted as under:


YearCash Flow

1 -5,00,00,000.00 Initial Investment

2 50,00,000.00 Regular Cash Flow

3 50,00,000.00 Regular Cash Flow

4 50,00,000.00 Regular Cash Flow

5 50,00,000.00 Regular Cash Flow

6 50,00,000.00 Regular Cash Flow

7 50,00,000.00 Regular Cash Flow

8 50,00,000.00 Regular Cash Flow

9 50,00,000.00 Regular Cash Flow

10 -1,50,00,000.00 Additional Investment

11 50,00,000.00 Regular Cash Flow

12 50,00,000.00 Regular Cash Flow

13 50,00,000.00 Regular Cash Flow

14 50,00,000.00 Regular Cash Flow

15 50,00,000.00 Regular Cash Flow

16 50,00,000.00 Regular Cash Flow

17 50,00,000.00 Regular Cash Flow

18 50,00,000.00 Regular Cash Flow

19 50,00,000.00 Regular Cash Flow

20 50,00,000.00 Regular Cash Flow

Internal Rate of Return= IRR (ALL THE CASH FLOW CELLS IN EXCEL) = 4.38%